An object moving at 49 m/s was able to increase its velocity to 128 m/s within 8.44 seconds. What is the acceleration of the object?
1
Expert's answer
2020-09-21T08:28:03-0400

We can find the acceleration of the object from the formula:


"a = \dfrac{v-v_0}{t} = \dfrac{128 \ \dfrac{m}{s}-49 \ \dfrac{m}{s}}{8.44 \ s} = 9.36 \ \dfrac{m}{s^2}."

Answer:

"a = 9.36 \ \dfrac{m}{s^2}."